OpenStudy (anonymous):

Working together, two pumps can drain a certain pool in 4 hours. If it takes the older pump 9 hours to drain the pool by itself, how long will it take the newer pump to drain the pool on its own?

OpenStudy (gorv):

1/9+1/x=1/4 (x+9)/9x=1/4 (x+9)*4=9x 4x+36=9x 5x=36 x=36/5